Find the length of one side of an equilateral triangle where each following side is 2x-7,x+y-9 and y+5

Answer by ikleyn(52780)

.

All three sides of an equilateral triangle have equal lengths.


It gives you these equations


    2x - 7 = y + 5,       (1)

    x + y - 9 = y + 5.    (2)


It is your system of 2 equations in 2 unknowns.


From equation (2),


    x = 5 + 9 = 14.


Substitute it into equation (1), replacing there x by 14


     2*14 - 7 =   y + 5.


It implies

     28 - 7 = y + 5

     21     = y + 5

      y     = 21 - 5 = 16.


Now, calculate the length of one side, for example, the length of the side 2x-7


     2x - 7 = 2*14 - 7 = 21.


Check that the other sides also have the length of 21


    x + y - 9 = 14 + 16 - 9 = 21

    y + 5 = 16 + 5 = 21.


ANSWER.  The side length of this equilateral triangle is 21 units.

Solved.